# TriMediQ: A Triplet-Structured Approach for Interactive Medical Question Answering

### v1

@@ Abstract
Large Language Models (LLMs) perform strongly in static and single-turn medical Question Answer (QA) benchmarks, yet such settings diverge from the iterative information gathering process required in practical clinical consultations. The MEDIQ framework addresses this mismatch by recasting the diagnosis as an interactive dialogue between a patient and an expert system, but the reliability of LLMs drops dramatically when forced to reason with dialogue logs, where clinical facts appear in sentences without clear links. To bridge this gap, we introduce TriMediQ, a triplet-structured approach that summarises patient responses into triplets and integrates them into a Knowledge Graph (KG), enabling multi-hop reasoning. We introduce a frozen triplet generator that extracts clinically relevant triplets, using prompts designed to ensure factual consistency. In parallel, a trainable projection module, comprising a graph encoder and a projector, captures relational information from the KG to enhance expert reasoning. TriMediQ operates in two steps: (i) the projection module fine-tuning with all LLM weights frozen; and (ii) using the fine-tuned module to guide multi-hop reasoning during inference. We evaluate TriMediQ on two interactive QA benchmarks, showing that it achieves up to 10.4% improvement in accuracy over five baselines on the iMedQA dataset. These results demonstrate that converting patient responses into structured triplet-based graphs enables more accurate clinical reasoning in multi-turn settings, providing a solution for the deployment of LLM-based medical assistants  [^1] The complete code and public datasets are available at: Anonymised Github..
@@ 1 Introduction
In this work, we propose TriMediQ, which is an approach that extends the MEDIQ framework while keeping the design of the patient and expert systems unchanged [21]. The patient system receives the complete patient record along with the current follow-up question, and returns a factual response by selecting only the directly relevant atomic facts from the record. The key difference is that TriMediQ provides the expert system with a triplet-structured KG constructed from UMLS-style triplets [2] such as (Patient, Has_Symptom, Fatigue). These triplets are incrementally accumulated during the interaction and organised into a semantic KG that captures both the factual content and the relationships among entities. Since effective reasoning requires understanding not only individual facts but also their interconnections, we employ a trainable graph encoder to capture relational structures and a projection module that maps the graph into a prefix representation for the expert, thereby enabling multi-hop reasoning.

### v2

# Triplet-Structured Knowledge Integration for Multi-Turn Medical Reasoning

@@ Abstract
Large Language Models (LLMs) have shown strong performance on static medical Question Answering (QA) tasks, yet their reasoning often deteriorates in multi-turn clinical dialogues where patient information is scattered across turns. This paper introduces TriMediQ, a triplet-structured approach that enhances the reasoning reliability of LLMs through explicit knowledge integration. TriMediQ first employs a frozen triplet extraction LLM to convert patient responses into clinically grounded triplets, ensuring factual precision via constrained prompting. These triplets are incorporated into a patient-specific Knowledge Graph (KG), from which a trainable projection module consisting of a graph encoder and a projector captures relational dependencies while keeping all LLM parameters frozen. During inference, the projection module guides multi-hop reasoning over the KG, enabling coherent clinical dialogue understanding. Experiments on two interactive medical QA benchmarks show that TriMediQ achieves up to 10.4% improvement in accuracy over five existing baselines on the iMedQA dataset. These results demonstrate that structuring patient information as triplets can effectively improve the reasoning capability of LLMs in multi-turn medical QA.
@@ 1 Introduction
In this work, we propose TriMediQ, which is an approach that extends the MEDIQ framework while retaining the design of the patient and expert systems [21]. The patient system receives the complete patient record along with the current follow-up question, and returns a factual response by selecting only the directly relevant atomic facts from the record. The key difference is that TriMediQ provides the expert system with a triplet-structured KG constructed from UMLS-style triplets [2] such as (Patient, Has_Symptom, Fatigue). These triplets are incrementally accumulated during the interaction and organised into a semantic KG that captures both the factual content and the relationships among entities. Since effective reasoning requires understanding not only individual facts but also their interconnections, we employ a trainable graph encoder to capture relational structures and a projection module that maps the graph into a prefix representation for the expert, thereby enabling multi-hop reasoning.